If a single electron in an excited hydrogen atom is occupying the 3rd energy level and then relaxes back to the ground state, how much energy is released in the form of electromagnetic radiation?
Answer #1

∆E =13.6×(1/n​​​​​​12 - 1/n22)

= 13.6×(1/12 -1/32)

= 13.6×8/9

= 12.1 ev

Energy released = 12.1 ev